At 20:10 UTC Andrea informed me about an unlock; the first attempt to relock failed, and he contacted me in case it was necessary to check the WI CH ( 67546 ).
Since I was awake, I connected to follow the situation live. While adjusting the WI CH power by rotating the waveplate (action completed at 20:28 UTC), we noticed that the flag TCS_NI_CO2_PWROUT was red.
Looking more carefully at the signals, I found a sudden drop (about a factor 2) in the NI CO2 laser power at 19:40 UTC, clearly visible in PWROUT, ISSIN, POWER_CH_PICKOFF. Interestingly, the effect was not as evident in PWRLAS and CO2laser_TE. See figure attached.
Since the change seemed real, I checked the NI CO2 laser powers individually (by closing flip mirrors) and restored the correct values injected into the ITF (action completed at 20:50 UTC).
I also asked Andrea to go to the TCS room to verify whether anything unusual could be seen or heard macroscopically, but no clear culprit was identified. The cause of the power drop remains to be further investigated.
Currently, the DMS flag TCS_NI_CO2_PWROUT will remain red, as I had to rotate the waveplate to send more power toward the ITF. The reference value should be re-centered, but we prefer to wait a few days before applying the change.